Required information [The following information applies to the questions displayed below.] The general ledger of Jackrabbit Rentals at January 1, 2021, includes the following account balances: Accounts Debits Credits Cash $ 46,500 Accounts Receivable 30,700 Land 115,800 Accounts Payable 15,800 Notes Payable (due in 2 years) 35,000 Common Stock 105,000 Retained Earnings 37,200 Totals $ 193,000 $ 193,000 The following is a summary of the transactions for the year: 1. January 12 Provide services to customers on account, $67,400. 2. February 25 Provide services to customers for cash, $77,800. 3. March 19 Collect on accounts receivable, $46,200. 4. April 30 Issue shares of common stock in exchange for $35,000 cash. 5. June 16 Purchase supplies on account, $13,100. 6. July 7 Pay on accounts payable, $11,800. 7. September 30 Pay salaries for employee work in the current year, $69,200. 8. November 22 Pay advertising for the current year, $23,000. 9. December 30 Pay $3,400 cash dividends to stockholders. The following information is available for the adjusting entries. Accrued interest on the notes payable at year-end amounted to $3,000 and will be paid January 1, 2022. Accrued salaries at year-end amounted to $2,000 and will be paid on January 5, 2022. Supplies remaining on hand at the end of the year equal $2,800.

7. Prepare an adjusted trial balance.
Answer #1

Entries

Account receivable      67400

To sales 67400

Cash      77800

To sales 77800

Cash 46200

To Account receivble 46200

Cash 35000

To Common Stock 35000

Purchases 13100

To Account Payable 13100

Account Payable 11800

Cash 11800

Salaries 69200

To Cash   69200

Advertising Expense 23000

To Cash 23000    

Dividend Paid 3400

To Cash 3400

Interest 3000

To Accrued Interest 3000

Salaries 2000

To Accrued Salaries 2000

Adjusted account Balances

Cash ( 46500 + 77800 + 46200 + 35000 – 11800 – 69200 – 23000 – 3400) = 98100

Account Receivable ( 30700 + 67400 – 46200 ) = 51900

Land = 115800

Account Payable = (15800 + 13100 – 11800) = 17100

Notes Payable = 35000

Common Stock = (105000 + 35000) = 1400001

Retained Earnings = 37,200

Sales = (67400 + 77800) = 145200

Purchases = 13100

Salaries = (69200 + 2000) = 71200

Advertising Expense = 23000

Dividend Paid = 3400

Accrued Salaries =2000

Accrued Interest = 3000

Interest = 3000

Trial Balance